Beta company makes and sells wadgets. Before the start of 2018, Alpha budgeted to produce and sell 24,000 wadgets. However, the company actually ended up producing and selling 27,400 wadgets in 2018. Beta budgeted to sell each wadget for $18, but the actual average selling price for each wadget was $21.50. For variable costs, beta budgeted that each wadget would use $3 of direct material costs, but the actual direct material cost was $2.50 per wadget. Please answer the following questions:

1. What will total sales be in the Flex Budget?

2. What will total direct material costs be in the Flex Budget?
Answer #1

Amount of Sales to be recognized in Flex Budget = Actual Sales Volume * Original Budgeted Selling Price

= 27400 * 18 $ = 493,200 $

Direct Material Cost to be recognized = Ratio of Actual Volume achieved budgeted Volume * Original Budget Variable Cost

= 27400 * 3 $ =  82,200 $

THANK YOU
HOPING FOR A POSITIVE RESPONSE